### v1

# Proxy-small objects present compactly generated categories

Abstract.
We develop a correspondence between presentations of compactly generated triangulated categories as localizations of derived categories of ring spectra and proxy-small objects, and explore some consequences. In addition, we give a characterization of proxy-smallness in terms of coproduct preservation of the associated corepresentable functor ‘up to base change’.
1. Introduction
Proxy-smallness is a relatively new concept that was introduced in [5, 6] as a weakening of smallness, also called compactness, in triangulated categories. It is general enough that the residue field of any noetherian local ring is proxy-small in the derived category, but strong enough to yield useful formulas for local cohomology. The most striking application has been in commutative algebra. Dwyer, Greenlees, and Iyengar proved [5]*Theorem 9.4 that if a commutative noetherian ring $R$ is complete intersection then every object of $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$ is proxy-small and they asked if this property characterized complete intersections. This remarkable statement turns out to be true, as was proved by Pollitz [16]. This opens many doors. By analogy it furnishes us with an elegant and homotopy invariant definition of complete intersection in derived settings, and it supports rather appealing, if speculative, connections to questions concerning finite generation of cohomology. The purpose of this article is to give a conceptual interpretation of proxy-smallness in analogy with the definition of compactness, that is in terms of some functor preserving coproducts.
Let $\mathsf{K}$ be an enhanced compactly generated triangulated category. Given a compact object $k\in\mathsf{K}$ there is an equivalence between the localizing subcategory $\operatorname{\mathrm{loc}}(k)$ generated by $k$ and the derived category $\mathsf{D}(E)$ where $E$ is the derived endomorphism ring $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(k,k)$ of $k$. This equivalence is given by the adjoint pair of functors $(-)\otimes_{E}k$ and $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(k,-)$, and boils down to the statements $E=\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(k,k)$ and $E\otimes_{E}k=k$, since both functors preserve coproducts. For a non-compact object $P$ this statement is false in general —the localizing subcategory $\operatorname{\mathrm{loc}}(P)$ depends on how $P$ sits in $\mathsf{K}$ in a subtle way, and not only on its endomorphism ring. For instance, in the derived category of abelian groups $\mathsf{D}(\mathbb{Z})$ the objects $\mathbb{Z}_{p}$ and $\mathbb{Z}[p^{-1}]/\mathbb{Z}$ have the same derived endomorphism ring but generate different localizing subcategories. The slogan is: giving a proxy-small object $P$ is the same thing as giving a presentation of the compactly generated localizing subcategory $\operatorname{\mathrm{loc}}(P)$ as a quotient of $\mathsf{D}(\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,P))$ (see 4.9).
This shows that proxy-smallness is ubiquitous. Moreover, as we point out, the proxy-small condition is hidden behind fundamental constructions such as passing from the derived category of a coherent ring $R$ to its ind-coherent derived category $\mathsf{K}(\operatorname{\mathsf{Inj}}R)$.
Along the way, in Theorem 4.3, we give a purely functorial characterization of the proxy-small condition when the ambient category satisfies the telescope conjecture. An object $P$ with $E=\operatorname{\mathbf{R}Hom}(P,P)$ is proxy-small if and only if $\operatorname{\mathbf{R}Hom}(P,-)\otimes_{E}P$ preserves coproducts. We also consider the relative version, where one enhanced triangulated category acts on another. The paper concludes with a discussion of the closure conditions the proxy-small objects satisfy as well as some speculations and applications.
2. Proxy-smallness revisited
We begin with some ‘recollections’ on proxy-smallness and effective constructibility, both in the sense of [6]. These are ‘recollections’ rather than recollections because the definition of proxy-smallness we use is slightly more general than the original and we give a different perspective, with some new results, on effective constructibility.
Let us fix a compactly generated stable $\infty$-category (i.e. an $\aleph_{0}$-presentable stable $\infty$-category) with homotopy category $\mathsf{K}$. Essentially $\mathsf{K}$ is a compactly generated triangulated category (see [12]*Remark 1.4.4.3) and we can speak of mapping spaces, homotopy (co)limits, and so on. Nothing is lost if the reader prefers to think of instead fixing a differential graded, or model, category with compactly generated homotopy category. In fact, essentially all arguments take place at the level of homotopy categories. We denote the compact objects $\mathsf{K}^{\mathrm{c}}$ and recall that this category is, by definition, essentially small.
Throughout by ‘localization’ we will mean left Bousfield localization, i.e. a functor with a fully faithful right adjoint. For a set of objects $\mathsf{X}$ of $\mathsf{K}$ we denote by $\operatorname{thick}(\mathsf{X})$ and $\operatorname{\mathrm{loc}}(\mathsf{X})$ the smallest triangulated subcategories containing $\mathsf{X}$ and closed under summands and arbitrary coproducts respectively.
Definition 2.1.
An object $P\in\mathsf{K}$ is a proxy-small object with small set of proxies $\mathbf{p}$ if
$\mathbf{p}\subseteq\operatorname{thick}(P)\cap\mathsf{K}^{\mathrm{c}}\text{ and }P\in\operatorname{\mathrm{loc}}(\mathbf{p}).$
We will often just say that $P$ is proxy-small without fixing a set of proxies.
Throughout we regard $\mathbf{p}$ as a full subcategory of $\mathsf{K}$. Using the enhancement we may consider the ind-completion $\mathsf{D}(\mathbf{p})$ of $\operatorname{thick}(\mathbf{p})$ and this identifies with the full subcategory $\operatorname{\mathrm{loc}}(\mathbf{p})$ of $\mathsf{K}$.
Remark 2.2.
If $P$ is proxy-small then $\operatorname{thick}(P)\cap\mathsf{K}^{\mathrm{c}}$ is a canonical set of proxies.
Remark 2.3.
If $P$ is proxy-small with set of proxies $\mathbf{p}$ then
$\operatorname{\mathrm{loc}}(\mathbf{p})=\operatorname{\mathrm{loc}}(P)=\operatorname{\mathrm{loc}}(\operatorname{thick}(P)\cap\mathsf{K}^{\mathrm{c}}).$
Remark 2.4.
This is a small generalization of the original definition of Dwyer, Greenlees, and Iyengar [6]. Instead of insisting there be a single small proxy we allow a set of such. This has the advantage that it does not depend on a choice of proxy if we just take $\mathbf{p}=\mathsf{K}^{\mathrm{c}}\cap\operatorname{thick}(P)$.
In fact one could, and in some cases should, go further: we could also replace $P$ by a set of objects. Everything works just as well if one takes this extra step. In our primary cases of interest we can get away with a single object $P$ and so we stick to this setting—replacing $P$ by a subcategory is essentially cosmetic.
Let us consider the ring spectrum $E=\operatorname{\mathbf{R}Hom}(P,P)$ and the full subcategory $\mathbf{p}$ of $\mathsf{K}$ and form the following diagram
[…]
consisting of the various tensor-hom adjunctions, and which commutes in the obvious ways by construction. Namely, the unlabelled left adjoints are given as follows: $\mathsf{D}(E)\longrightarrow\mathsf{K}$ is $(-)\otimes_{E}P$ which factors via $\operatorname{\mathrm{loc}}(P)=\operatorname{\mathrm{loc}}(\mathbf{p})$ by virtue of preserving coproducts, $\mathsf{D}(\mathbf{p})\longrightarrow\operatorname{\mathrm{loc}}(\mathbf{p})$ is the equivalence determined by the universal property of ind-completion (that is, by sending $\operatorname{\mathbf{R}Hom}_{\mathbf{p}}(-,p)$ to $p$ for $p\in\mathbf{p}$), and $\mathsf{D}(\mathbf{p})\longrightarrow\mathsf{K}$, which we denote by $(-)\otimes_{\mathbf{p}}\mathbf{p}$, is also determined by the universal property.
Remark 2.6.
In most cases of interest it is possible to pick a single proxy and take $\mathbf{p}=\{p\}$. In this case we can just deal with $B=\operatorname{\mathbf{R}Hom}(p,p)$, for instance $\mathsf{D}(\mathbf{p})\cong\mathsf{D}(B)$ and $\mathsf{D}(B)\longrightarrow\mathsf{K}$ is the functor $(-)\otimes_{B}p$ which factors via $\operatorname{\mathrm{loc}}(p)$. The reader may find it helpful to keep this case in mind.
For the rest of this section $P$ is a proxy-small object in $\mathsf{K}$ with a set of proxies $\mathbf{p}$. We are primarily concerned with understanding
$\Gamma=i_{*}i^{!}\cong\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},-)\otimes_{\mathbf{p}}\mathbf{p}.$
To this end let us begin by understanding the composite $\mathsf{D}(E)\longrightarrow\mathsf{D}(\mathbf{p})$.
Lemma 2.7.
The functor $\mathsf{D}(E)\longrightarrow\mathsf{D}(\mathbf{p})$ is naturally isomorphic to
$-\otimes_{E}\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},P)$
where $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},P)$ is the $E^{\mathrm{op}}\otimes\mathbf{p}$-module given by restricting $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(-,-)$ to the indicated objects.
Proof.
The composite $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},-\otimes_{E}P):\mathsf{D}(E)\longrightarrow\mathsf{D}(\mathbf{p})$ preserves coproducts since $\mathbf{p}$ consists of compacts. As we vary $M\in\mathsf{D}(E)$, by tracing the image of the identity under the chain of isomorphisms
$\operatorname{\mathbf{R}Hom}_{K}(M\otimes_{E}P,M\otimes_{E}P)$ $\cong\operatorname{\mathbf{R}Hom}_{E}(M,\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,M\otimes_{E}P))$
$\cong\operatorname{\mathbf{R}Hom}_{E}(M,\operatorname{\mathbf{R}Hom}_{\mathbf{p}}(\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},P),\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},M\otimes_{E}P)))$
$\cong\operatorname{\mathbf{R}Hom}_{\mathbf{p}}(M\otimes_{E}\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},P),\operatorname{\mathbf{R}Hom}_{E}(\mathbf{p},M\otimes_{E}P)),$
where the third isomorphism is fully faithfulness of $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},-)$ on $\operatorname{\mathrm{loc}}(\mathbf{p})$ and the others are adjunction isomorphisms, we obtain a natural map
$-\otimes_{E}\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},P)\longrightarrow\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},-\otimes_{E}P).$
This is evidently an isomorphism at $E$, and both functors preserve coproducts, so this natural transformation is an isomorphism as claimed. ∎
Remark 2.8.
It follows that the right adjoint functor $\mathsf{D}(\mathbf{p})\longrightarrow\mathsf{D}(E)$ is given by
$\operatorname{\mathbf{R}Hom}_{\mathbf{p}}(\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},P),-).$
We now prove the key fact concerning proxy-smallness.
Theorem 2.9.
Assume $P$ is proxy-small. The functor $(-)\otimes_{E}P\colon\mathsf{D}(E)\longrightarrow\operatorname{\mathrm{loc}}(P)$ is a localization. Moreover, it preserves products.
Proof.
Our starting point is the observation that $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)$ induces an equivalence from $\operatorname{thick}(P)$ to $\mathsf{D}(E)^{\mathrm{c}}$ with inverse $(-)\otimes_{E}P$. In particular, $\mathbf{p}\subseteq\operatorname{thick}(P)$ embeds fully faithfully into $\mathsf{D}(E)$ which induces a colimit and compact preserving embedding $\lambda\colon\operatorname{\mathrm{loc}}(\mathbf{p})\longrightarrow\mathsf{D}(E)$. While the localizing subcategory $\operatorname{\mathrm{loc}}(P)$ must be taken in $\mathsf{K}$, we note that $\operatorname{\mathrm{loc}}(\mathbf{p})$ is unambiguous here: $\mathbf{p}$ consists of objects that are compact in both $\mathsf{D}(E)$ and $\mathsf{K}$, and so it does not matter where we form the localizing subcategory $\mathbf{p}$ generates. By Brown representability $\lambda$ has a right adjoint $\rho$. Of course, this is overkill: the right adjoint is just the obvious restriction functor (up to identifying $\operatorname{\mathrm{loc}}(\mathbf{p})$ and $\mathsf{D}(\mathbf{p})$).
We claim that $\rho\cong(-)\otimes_{E}P$. Since both of these functors preserve colimits they are determined by their value at $E$. We use the identification of $\operatorname{thick}(P)$, where the thick closure is taken in $\operatorname{\mathrm{loc}}(\mathbf{p}),$ with $\mathsf{D}(E)^{\mathrm{c}}$ again. The latter contains $\lambda(\mathbf{p})$ and so
$\rho(E)=\operatorname{\mathbf{R}Hom}_{\mathsf{D}(E)}(-,E)|_{\lambda(\mathbf{p})}\cong\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(-,P)|_{\mathbf{p}}.$
Now the representable functor on the right corresponds to $P\in\operatorname{\mathrm{loc}}(P)=\operatorname{\mathrm{loc}}(\mathbf{p})$ under the equivalence $\mathsf{D}(\mathbf{p})\cong\operatorname{\mathrm{loc}}(\mathbf{p})$, and we can conclude that $\rho\cong(-)\otimes_{E}P$. It follows that the right adjoint $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)$ of $(-)\otimes_{E}P$ is fully faithful, because the left adjoint $\lambda$ of $\rho$ is so, showing that the latter is a localization. Being a right adjoint $\rho\cong(-)\otimes_{E}P$ preserves products as claimed. ∎
Remark 2.10.
The theorem gives a rather pleasant perspective on proxy-smallness: Given an object $X$ of $\mathsf{K}$ with derived endomorphism ring $E$, there is always an adjunction
${\mathsf{D}(E)}$${\operatorname{\mathrm{loc}}(X)}$
where $\mathsf{D}(R)\longrightarrow\operatorname{\mathrm{loc}}(X)$ is given by $(-)\otimes_{R}X$. However, there is no reason for this to be a localization. Proxy-smallness of $X$ guarantees that it is.
This alone is already interesting in a number of cases.
Corollary 2.11.
Let $(R,\mathfrak{m},k)$ be a commutative noetherian local ring and let $E$ denote the derived endomorphism ring of $k$. Then $\mathsf{D}(E)\longrightarrow\mathsf{D}_{\{\mathfrak{m}\}}(R)=\operatorname{\mathrm{loc}}(k)$ is a localization.
Proof.
The residue field $k$ is always proxy-small, with proxy given by, for example, the Koszul complex on $\mathfrak{m}$. ∎
Example 2.12.
Consider $R=k[x]/(x^{2})$. In this case, $E=k[\theta]$ with $\theta$ in degree $1$ and no higher homotopical structure (i.e. $E$ is formal). We have
$\mathsf{D}^{\mathrm{b}}(R)=\operatorname{thick}(k)\subseteq\operatorname{\mathrm{loc}}(k)=\mathsf{D}(R)$
and a fully faithful embedding of $\mathsf{D}(R)$ into $\mathsf{K}(\operatorname{\mathsf{Inj}}R)$ the ind-coherent sheaves on $\operatorname{Spec}R$ given by taking $K$-injective resolutions. The compact objects of $\mathsf{K}(\operatorname{\mathsf{Inj}}R)$ are the image of $\mathsf{D}^{\mathrm{b}}(R)$ under this embedding and so an injective resolution of $k$ is a compact generator. Thus there is an equivalence of $\mathsf{K}(\operatorname{\mathsf{Inj}}R)$ with $\mathsf{D}(E)$ which identifies the above embedding with $\operatorname{\mathbf{R}Hom}(k,-)\colon\mathsf{D}(R)\longrightarrow\mathsf{D}(E)$. The left adjoint to this embedding is the localization $\mathsf{D}(E)\longrightarrow\mathsf{D}(R)=\mathsf{D}_{\{\mathfrak{m}\}}(R)$.
This extends to any artinian commutative local ring, and in fact it is more general still.
Example 2.13.
Let $R$ be a coherent ring and $G$ a classical generator for $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$. We note that such a $G$ does not exist for all rings, not even commutative ones. Then $G$ is certainly proxy-small: $\operatorname{thick}(G)=\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$ which contains all perfect complexes and it follows that $\operatorname{\mathrm{loc}}(G)=\mathsf{D}(R)$ which is generated by the perfect complexes.
Hence, letting $E$ denote the derived endomorphism ring of $G$, we see $\mathsf{D}(R)$ is a localization of $\mathsf{D}(E)$ and $\mathsf{D}(E)^{\mathrm{c}}\cong\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$. Thus $\mathsf{D}(E)\cong\mathsf{K}(\operatorname{\mathsf{Inj}}R)$ and this localization is the usual one up to equivalence.
Remark 2.14.
Allowing $P$ to consist of a set of objects, we can take $P=\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$ for any coherent ring $R$ (or scheme) and the result can be identified with Krause’s recollement [11] (as extended to coherent rings by Šťovíček [19]) connecting $\mathsf{K}(\operatorname{\mathsf{Inj}}R)$ and $\mathsf{D}(R)$ essentially for free from the formalism of proxy-smallness. We are grateful to Šťovíček for pointing this out.
Using Theorem 2.9 we can give an efficient proof of effective constructibility as in [6]*Theorem 4.10.
Proposition 2.15.
There is a natural isomorphism
$\Gamma\cong\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P.$
Proof.
We know, since $\mathbf{p}$ consists of compact objects, there is an isomorphism $\Gamma\cong\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},-)\otimes_{\mathbf{p}}\mathbf{p}$. Hence, for $X\in\mathsf{K}$ we have
[…]
where the top horizontal isomorphism comes from factoring $\mathsf{D}(E)\longrightarrow\mathsf{K}$ via $\mathsf{D}(\mathbf{p})$ and applying Lemma 2.7. So we just need to check that the vertical map is an isomorphism, and of course it suffices to show that
$\varepsilon\colon\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,X)\otimes_{E}\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},P)\longrightarrow\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},X)$
is an isomorphism. One can see this directly from chasing diagram (2.5) using Lemma 2.7 and Theorem 2.9. ∎
Example 2.16.
As the reader may have already divined from the setup it is not necessarily the case that $\mathsf{D}(E)\longrightarrow\mathsf{K}$ is fully faithful. In fact, it is not even necessarily conservative.
Consider $\mathsf{K}=\mathsf{D}(\mathbb{Z})$ and let $P$ be the injective envelope of $\mathbb{F}_{p}$. Then $E=\operatorname{\mathbf{R}Hom}_{\mathbb{Z}}(P,P)$ is concentrated in degree 00 and isomorphic to the $p$-adic integers $\mathbb{Z}_{p}$. Note that $P$ is proxy-small (e.g. with small proxy $\mathbb{F}_{p}$ cf. Example 6.3). One can easily check that $\mathbb{Q}_{p}$, the function field of $\mathbb{Z}_{p}$, is killed by
$-\otimes_{\mathbb{Z}_{p}}P\colon\mathsf{D}(\mathbb{Z}_{p})\longrightarrow\mathsf{D}(\mathbb{Z}).$
This is not surprising when we keep in mind Theorem 2.9 that tells us $\mathsf{D}(E)\longrightarrow\mathsf{K}$ factors via a localization (which has no reason to be the trivial one).
3. Detecting smashing localizations
We continue the notation of the previous sections, with the caveat that we no longer assume that $P$ is necessarily proxy-small. If $P$ is proxy-small then the functor $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts, even though $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)$ does not if $P$ is not compact. It is then natural to ask what we can deduce about a general $P$ with the property that $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts.
Set $F=-\otimes_{E}P$ and $G=\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)$ for brevity, and assume that $FG$ preserves coproducts. We consider these functors, as in diagram (2.5), as an adjunction relating $\mathsf{D}(E)$ and $\mathsf{K}$. The counit of this adjunction is denoted $\varepsilon\colon FG\longrightarrow\operatorname{id}_{\mathsf{K}}$. The following lemma is implicit in said diagram, but let us make it explicit.
Lemma 3.1.
The essential image of $FG$, denoted $\operatorname{im}FG$, is contained in $\operatorname{\mathrm{loc}}(P)$.
Proof.
It is clear that $\operatorname{im}FG\subseteq\operatorname{im}F$. We have $FE=P$ and $F$ preserves coproducts and is exact so $\operatorname{im}F\subseteq\operatorname{\mathrm{loc}}(P)$ by the usual arguments. ∎
Now let us use that $FG$ preserves coproducts.
Lemma 3.2.
The map $\varepsilon_{FG}\colon(FG)^{2}\longrightarrow FG$ is an isomorphism.
Proof.
Since $FG$ preserves coproducts
$\mathsf{N}=\{X\in\mathsf{K}\mid FG(X)\stackrel{{\scriptstyle\varepsilon}}{{\longrightarrow}}X\text{ is an isomorphism}\}$
is a localizing subcategory. On $P$ we get
$\varepsilon_{P}\colon FG(P)=\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,P)\otimes_{E}P\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}P$
and so $\operatorname{\mathrm{loc}}(P)\subseteq\mathsf{N}$. The statement then follows from Lemma 3.1. ∎
Since $\mathsf{K}$ is compactly generated $\operatorname{\mathrm{loc}}(P)$ is well-generated [15, Theorem 4.4.9] and so by Brown representability [15, Theorem 8.3.3] there is for all $X$ a localization triangle
$\Gamma_{P}X\longrightarrow X\longrightarrow L_{P}X$ (3.3)
with $\Gamma_{P}X$ in $\operatorname{\mathrm{loc}}(P)$ and $L_{P}X$ in $\operatorname{\mathrm{loc}}(P)^{\perp}$.
Lemma 3.4.
The natural map $FG(\varepsilon)\colon(FG)^{2}\longrightarrow FG$ is an isomorphism.
Proof.
Since $FG$ preserves coproducts
$\mathsf{N}=\{X\in\mathsf{K}\mid FG(\varepsilon)\colon(FG)^{2}(X)\longrightarrow FG(X)\text{ is an isomorphism}\}$
is a localizing subcategory. The map $FG(\varepsilon)$ is an isomorphism on $P$ (in fact the same one as in Lemma 3.2) and hence on all of $\operatorname{\mathrm{loc}}(P)$. On the other hand, if $X\in\operatorname{\mathrm{loc}}(P)^{\perp}$ then $G(X)=0$ by definition and so $FG(X)=0$. Thus $\varepsilon_{X}=0$ and so $FG(\varepsilon_{X})\colon 0\longrightarrow 0$ is an isomorphism and we see $\operatorname{\mathrm{loc}}(P)^{\perp}\subseteq\mathsf{N}$. Thus $\mathsf{N}$ contains $\operatorname{\mathrm{loc}}(P)$ and $\operatorname{\mathrm{loc}}(P)^{\perp}$, from which we conclude using the localization triangle (3.3) that $\sf N=\mathsf{K}$. ∎
To summarize: we have shown that the coproduct preserving functor
$FG=\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P,$
equipped with the counit $\varepsilon\colon FG\longrightarrow\operatorname{id}_{\mathsf{K}}$, has essential image $\operatorname{\mathrm{loc}}(P)$ and that $FG(\varepsilon)$ and $\varepsilon_{FG}$ are isomorphisms. Thus provided we can show that $FG(\varepsilon)=\varepsilon_{FG}$ we would see that $FG$ is an acyclization functor and $\operatorname{\mathrm{loc}}(P)$ is smashing.
This is true on objects of $\operatorname{\mathrm{loc}}(P)$. Indeed, the naturality square at an object $X$ for $\varepsilon$ with respect to itself
${(FG)^{2}(X)}$${FG(X)}$${FG(X)}$${X}$$\scriptstyle{FG(\varepsilon_{X})}$$\scriptstyle{\varepsilon_{FG(X)}}$$\scriptstyle{\varepsilon_{X}}$$\scriptstyle{\varepsilon_{X}}$
shows these two transformations agree up to $\varepsilon_{X}$. In particular, if $X\in\operatorname{\mathrm{loc}}(P)$ then $\varepsilon_{X}$ is an isomorphism (e.g. by the proof of Lemma 3.2) and so we can cancel it to deduce that $FG(\varepsilon_{X})=\varepsilon_{FGX}$.
Applying $(FG)^{2}\longrightarrow FG$ to the localization triangle (3.3), and noting that $G(L_{P}X)=0$, we get a pair of maps of triangles
[…]
from which we see $FG(\varepsilon_{X})=\varepsilon_{FGX}$. Thus we have proved the following theorem.
Theorem 3.5.
Let $P$ be an object of $\mathsf{K}$, with derived endomorphism ring $E$, such that $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts. Then $\operatorname{\mathrm{loc}}(P)$ is smashing with acyclization functor $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$.∎
4. Characterization of proxy-smallness
We use Theorem 3.5 to give a characterization of proxy-small objects. Before proceeding let us give a quick reminder. We have an object $P$ of $\mathsf{K}$ with derived endomorphism ring $E$. This data induces a composite functor
$\mathsf{D}(E)\longrightarrow\operatorname{\mathrm{loc}}(P)\longrightarrow\mathsf{K}$
with a right adjoint, and we have observed $\mathsf{D}(E)\longrightarrow\operatorname{\mathrm{loc}}(P)$ is not necessarily an equivalence. However, it does induce an equivalence $\mathsf{D}(E)^{\mathrm{c}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\operatorname{thick}(P)$ between the perfect $E$-modules and the thick subcategory generated by $P$.
We have the following ‘boundedness’ result.
Lemma 4.1.
Let $P$ be an object of $\mathsf{K}$ with derived endomorphism ring $E$. Suppose that $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts. Then $\operatorname{\mathrm{loc}}(P)\cap\mathsf{K}^{\mathrm{c}}\subseteq\operatorname{thick}(P)$.
Proof.
Let $x$ be an object of $\operatorname{\mathrm{loc}}(P)\cap\mathsf{K}^{\mathrm{c}}$. The object $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,x)$ of $\mathsf{D}(E)$ can be written as a filtered homotopy colimit $\operatorname{hocolim}_{i}F_{i}$ where each $F_{i}$ lies in $\operatorname{thick}(E)$. The left adjoint $(-)\otimes_{E}P$ preserves homotopy colimits and so we get isomorphisms
$x\cong\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,x)\otimes_{E}P\cong(\operatorname{hocolim}_{i}F_{i})\otimes_{E}P\cong\operatorname{hocolim}_{i}(F_{i}\otimes_{E}P)$
the first via Theorem 3.5. Because $x$ is compact in $\operatorname{\mathrm{loc}}(P)$ the isomorphism from $x$ to $\operatorname{hocolim}_{i}(F_{i}\otimes_{E}P)$ factors through some $F_{i}\otimes_{E}P$. Hence $x$ is a retract of the object $F_{i}\otimes_{E}P\in\operatorname{thick}(P)$ and so $x$ lies in $\operatorname{thick}(P)$. ∎
Remark 4.2.
For $P$ as above we know $\operatorname{\mathrm{loc}}(P)$ is smashing, by Theorem 3.5, and so the inclusion functor $\operatorname{\mathrm{loc}}(P)\longrightarrow\mathsf{K}$ sends compacts to compacts. Thus, being compact in $\operatorname{\mathrm{loc}}(P)$ is the same as being compact in $\mathsf{K}$ and we do not need to distinguish these notions.
All this leads to the following characterization of proxy-smallness.
Theorem 4.3.
Let $P$ be an object of $\mathsf{K}$, with derived endomorphism ring $E$. Then $P$ is proxy-small if and only if the following two conditions hold:
  • (i)
    $\operatorname{\mathrm{loc}}(P)=\operatorname{\mathrm{loc}}(\operatorname{\mathrm{loc}}(P)\cap\mathsf{K}^{\mathrm{c}})$;
  • (ii)
    $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts.
Proof.
If $P$ is proxy-small then (i) holds by definition and (ii) holds by effective constructibility as in Proposition 2.15.
Suppose, on the other hand, the object $P$ verifies (i) and (ii). By Theorem 3.5 condition (ii) guarantees that $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ is the acyclization functor for $\operatorname{\mathrm{loc}}(P)$. By Lemma 4.1 if $x\in\operatorname{\mathrm{loc}}(P)$ is compact in $\mathsf{K}$ then $x$ already lies in $\operatorname{thick}(P)$. Condition (i) then ensures there are enough such compacts. ∎
Remark 4.4.
It was already shown in [6], provided $\operatorname{\mathrm{loc}}(P)$ has a compact generator, that proxy-smallness of $P$ implies (i) and (ii).
Corollary 4.5.
Suppose that $\mathsf{K}$ satisfies the telescope conjecture, i.e. every smashing subcategory of $\mathsf{K}$ is generated by the compact objects it contains. Then $P$ is proxy-small if and only if $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts.
Proof.
If the telescope conjecture holds for $\mathsf{K}$ then condition (ii) already implies condition (i) by Theorem 3.5. ∎
Corollary 4.6.
Let $\mathsf{J}$ be a localizing subcategory of $\mathsf{K}$ satisfying $\mathsf{J}=\operatorname{\mathrm{loc}}(\mathsf{J}\cap\mathsf{K}^{\mathrm{c}})$. If $E$ is a ring spectrum and $\pi\colon\mathsf{D}(E)\longrightarrow\mathsf{J}$ is a localization then $\pi(E)$ is proxy-small in $\mathsf{K}$.
Proof.
Being a left adjoint $\pi$ preserves coproducts, so it necessarily of the form $-\otimes_{E}\pi(E)$, with right adjoint given by $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\pi(E),-)$. By the assumption that $\pi$ is a localization we have a natural isomorphism $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\pi(E),-)\otimes_{E}\pi(E)\cong\operatorname{id}_{\mathsf{J}}$. Since the inclusion $i_{*}\colon\mathsf{J}\longrightarrow\mathsf{K}$ preserves compacts, it has a coproduct preserving right adjoint $i^{!}$. Then $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\pi(E),-)\otimes_{E}\pi(E)=i_{*}(\operatorname{\mathbf{R}Hom}_{\mathsf{J}}(\pi(E),i^{!}(-))\otimes_{E}\pi(E))=i_{*}i^{!}$ evidently preserves coproducts, and $\pi(E)$ is proxy-small by Theorem 4.3. ∎
Remark 4.7.
One can also prove the theorem by checking the definition of proxy-smallness directly using that $\mathsf{J}=\operatorname{\mathrm{loc}}(\pi(E))$, because $\pi$ is a localization, and applying Lemma 4.1.
Remark 4.8.
Choosing a set of small proxies $\mathbf{p}$ is the additional data of fixing another presentation of $\mathsf{J}$ as a derived category, namely $\mathsf{J}\cong\mathsf{D}(\mathbf{p})$.
Slogan 4.9.
Proxy-small objects are presentations of localizing subcategories generated by compacts: giving a proxy-small object $P$ is equivalent to giving a thick subcategory $\mathsf{R}$ of the compact objects $\mathsf{K}^{\mathrm{c}}$, a ring spectrum $E$, and a localization $\mathsf{D}(E)\longrightarrow\operatorname{\mathrm{loc}}(\mathsf{R})$.
Let us justify this. We saw already, via Theorem 2.9, that given a proxy-small $P$ with derived endomorphism ring $E$ we get a localization $\mathsf{D}(E)\longrightarrow\operatorname{\mathrm{loc}}(P)$ and (by definition) $\operatorname{\mathrm{loc}}(P)$ is generated by objects of $\mathsf{K}^{\mathrm{c}}$. Going in the other direction is precisely Corollary 4.6.
Remark 4.10.
In fact the localization corresponding to a proxy-small object also preserves products. This was shown in 2.9.
When $\mathsf{K}$ has a compact generator each compactly generated localizing subcategory comes with a canonical presentation. Let $\mathbf{p}$ be a set of compact objects of $\mathsf{K}$ and set
$\mathsf{J}^{\mathrm{c}}=\operatorname{thick}(\mathbf{p})\subseteq\operatorname{\mathrm{loc}}(\mathbf{p})=\mathsf{J}.$
We continue to denote by $i_{*}$ and $i^{!}$ the fully faithful inclusion $\mathsf{J}\longrightarrow\mathsf{K}$ and its right adjoint which is a localization. This gives rise to the acyclization functor $\Gamma=i_{*}i^{!}$ as before.
Proposition 4.11.
Suppose that $\mathsf{K}$ has a compact generator $g$. The object $\Gamma g$ is proxy-small. Moreover, one can take $\mathbf{p}$ as a set of small proxies.
Proof.
By assumption we have $\operatorname{thick}(g)=\mathsf{K}^{\mathrm{c}}\supseteq\mathsf{J}^{\mathrm{c}}$. Applying $\Gamma$ we obtain
$\operatorname{thick}(\Gamma g)\supseteq\Gamma\operatorname{thick}(g)=\Gamma(\mathsf{K}^{\mathrm{c}})\supseteq\Gamma(\mathsf{J}^{\mathrm{c}})=\mathsf{J}^{\mathrm{c}}\supseteq\mathbf{p}.$
By construction $\operatorname{\mathrm{loc}}(\Gamma g)=\mathsf{J}=\operatorname{\mathrm{loc}}(\mathbf{p})$ and so $\Gamma g$ is proxy-small, with set of proxies $\mathbf{p}$. ∎
Remark 4.12.
Even if $\mathsf{K}$ does not have a generator we can, using the expanded definition of proxy-smallness of Remark 2.1, consider $\Gamma\mathsf{K}^{\mathrm{c}}$ with set of small proxies $\mathsf{K}^{\mathrm{c}}$.
This gives the claimed canonical presentation which is, it seems, new as well as being rather striking.
Corollary 4.13.
Suppose that $\mathsf{K}$ has a compact generator $g$. Then $\mathsf{J}$ is a localization of $\mathsf{D}(\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\Gamma g,\Gamma g))$.∎
It is interesting to understand this phenomenon more thoroughly. In analogy with the situation in algebraic geometry one should perhaps think of this as some form of completion.
5. The relative setting
In this section we discuss a version of proxy-smallness in the setting of a closed monoidal triangulated category acting on another triangulated category. Our main motivation is to capture phenomena in algebraic geometry and representation theory where typically one needs additional closure conditions, coming from the action, to realize the connection between supports and thick subcategories.
Let $\mathsf{T}$ be a monoidal $\aleph_{0}$-presentable stable $\infty$-category with an associative and unital left action $\mathsf{T}\times\mathsf{K}\stackrel{{\scriptstyle\ast}}{{\longrightarrow}}\mathsf{K}$ on the $\aleph_{0}$-presentable stable $\infty$-category $\mathsf{K}$. We assume that the monoidal structure and action preserve colimits in each variable. In particular, at the level of homotopy categories we have an action of a monoidal triangulated category in the sense of [17] (except without the insistence on symmetry of the monoidal structure), and as before it is sufficient for the reader to just have this in mind. A prototypical example is the action of $\mathsf{T}$ on itself via the monoidal structure.
Given a full subcategory $\mathsf{A}$ of $\mathsf{K}$ we denote by $\operatorname{\mathrm{loc}}^{\ast}(\mathsf{A})$ and $\operatorname{thick}^{\ast}(\mathsf{A})$ the smallest localizing $\mathsf{T}$-submodule of $\mathsf{K}$ containing $\mathsf{A}$ and the smallest thick $\mathsf{T}^{\mathrm{c}}$-submodule of $\mathsf{K}$ containing $\mathsf{A}$ respectively. These are given by closing under the action, e.g. taking $\{X\ast A\mid X\in\mathsf{T},A\in\mathsf{A}\}$ and then taking the smallest localizing subcategory containing these objects gives $\operatorname{\mathrm{loc}}^{\ast}(\mathsf{A})$.
Definition 5.1.
An object $P\in\mathsf{K}$ is a $\ast$-proxy-small object with small set of proxies $\mathbf{p}$ if
$\mathbf{p}\subseteq\mathsf{K}^{\mathrm{c}}\cap\operatorname{thick}^{\ast}(P)\text{ and }P\in\operatorname{\mathrm{loc}}^{\ast}(\mathbf{p}).$
As before, we will often just say that $P$ is $\ast$-proxy-small without fixing a set of proxies.
Remark 5.2.
The above definition is equivalent to asking that $\mathsf{T}^{\mathrm{c}}\ast P=\{x\ast P\mid x\in\mathsf{T}^{\mathrm{c}}\}$ is proxy-small in the absolute sense.
Remark 5.3.
As in the absolute setting it makes perfect sense to replace the object $P$ by a full subcategory and this only results in cosmetic changes.
This again gives rise to presentations of $\operatorname{\mathrm{loc}}^{\ast}(P)$, that are now $\mathsf{T}$-linear in an appropriate sense. Recall that $\mathsf{K}$ satisfies the $\ast$-telescope conjecture if every smashing $\ast$-submodule of $\mathsf{K}$ is generated by objects of $\mathsf{K}^{\mathrm{c}}$.
Theorem 5.4.
Let $P$ be an object of $\mathsf{K}$, let $\mathsf{G}$ be a set of compact generators for $\mathsf{T}$, and set $\mathsf{G}\ast P=\{g\ast P\mid g\in\mathsf{G}\}$. Let us moreover assume that $\mathsf{K}$ satisfies the $\ast$-telescope conjecture. Then the following statements are equivalent:
  • (i)
    $P$ is $\ast$-proxy-small;
  • (ii)
    the functor $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathsf{G}\ast P,-)\otimes_{\mathsf{G}\ast P}\mathsf{G}\ast P$ preserves coproducts;
  • (iii)
    there is a product preserving $\mathsf{T}$-linear localization $\mathsf{D}(\mathsf{G}\ast P)\longrightarrow\operatorname{\mathrm{loc}}^{\ast}(P)$ sending $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(-,P)$ to $P$.
Proof.
The equivalence of (i) and (ii) is a direct application of Corollary 4.5.
Now let us show that (i) and (iii) are also equivalent. Let $P$ be an $\ast$-proxy-small object of $\mathsf{K}$ and fix $\mathsf{G}$ as in the statement. Because $\operatorname{thick}(\mathsf{G})=\mathsf{T}^{\mathrm{c}}$ we know $\operatorname{thick}(\mathsf{G}\ast P)=\operatorname{thick}^{\ast}(P)$ (cf. [17]*Lemma 3.11) and so $\mathsf{G}\ast P$ is proxy-small in the absolute sense. Theorem 2.9 then furnishes us with the required localization and it only remains to construct the action of $\mathsf{T}$ on $\mathsf{D}(\mathsf{G}\ast P)$ and check linearity.
The compact objects of $\mathsf{D}(\mathsf{G}\ast P)$ are equivalent to $\operatorname{thick}(\mathsf{G}\ast P)=\operatorname{thick}^{\ast}(P)$ and so are canonically endowed with an action of $\mathsf{T}^{\mathrm{c}}$. Passing to ind-completions gives the desired action of $\mathsf{T}$ on $\mathsf{D}(\mathsf{G}\ast P)$ and linearity of the localization $\mathsf{D}(\mathsf{G}\ast P)\longrightarrow\operatorname{\mathrm{loc}}^{\ast}(P)$ then follows from the fact it is linear on generators and everything in sight preserving colimits.
Finally, suppose we are given a localization $\pi\colon\mathsf{D}(\mathsf{G}\ast P)\longrightarrow\operatorname{\mathrm{loc}}^{\ast}(P)$ as in (iii) of the statement. Then by $\mathsf{T}$-linearity $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(-,g\ast P)$ is sent to $g\ast P$, and by Corollary 4.6 the subcategory $\mathsf{G}\ast P$ of $\mathsf{K}$ is proxy-small. Using again that $\operatorname{thick}(\mathsf{G}\ast P)=\operatorname{thick}^{\ast}(P)$ we see $P$ is $\ast$-proxy-small. ∎
Remark 5.5.
The canonical choice is $\mathsf{G}=\mathsf{T}^{\mathrm{c}}$. Different choices satisfying $\mathsf{G}\subseteq\mathsf{G}^{\prime}$ give presentations, as in (iii), related by $\mathsf{T}$-linear localizations.
Example 5.6.
Let $T$ be a separated noetherian regular scheme of finite Krull dimension, $\mathcal{E}$ a vector bundle on $T$, $t\in\Gamma(T,\mathcal{E})$ a regular section, and let $X$ be the zero scheme of $t$. We take $\mathsf{T}=\mathsf{K}=\mathsf{D}(\operatorname{QCoh}X)$ the unbounded derived category of quasi-coherent sheaves which acts on itself via the derived tensor product. Then [18]*Lemma 4.1 asserts that every object $\mathcal{F}\in\mathsf{D}^{\mathrm{b}}(\operatorname{Coh}X)$ is $\otimes$-proxy-small.
Example 5.7.
Let $G$ be a finite group and let $k$ be a field whose characteristic divides the order of $G$. We take $\mathsf{T}=\mathsf{K}=\mathsf{D}(kG)$ acting on itself via the tensor product. Then every object of $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}kG)$ is $\otimes$-proxy-small.
6. Closure conditions
We give a brief discussion of the closure conditions the proxy-small objects satisfy. They are rather different than what one generally hopes for in the stable setting. We begin with the observation that proxy-smallness is preserved under taking coproducts.
Lemma 6.1.
The class of proxy-small objects is closed under suspensions and arbitrary co-products.
Proof.
It is clear that a suspension of a proxy-small object is proxy-small.
Let $\{P_{\lambda}\mid\lambda\in\Lambda\}$ be a set of proxy-small objects. Set $P=\oplus_{\lambda}P_{\lambda}$ and $\mathbf{p}_{\lambda}=\operatorname{\mathrm{loc}}(P_{\lambda})\cap\mathsf{K}^{\mathrm{c}}$. By virtue of being summands of $P$ each $P_{\lambda}$ is in $\operatorname{thick}(P)$ and hence $\mathbf{p}_{\lambda}\subseteq\operatorname{thick}(P)$ for each $\lambda$. On the other hand, $\operatorname{\mathrm{loc}}(P)=\operatorname{\mathrm{loc}}(P_{\lambda}\mid\lambda\in\Lambda)$ and $\operatorname{\mathrm{loc}}(P_{\lambda})=\operatorname{\mathrm{loc}}(\mathbf{p}_{\lambda})$ so
$\operatorname{\mathrm{loc}}(P_{\lambda}\mid\lambda\in\Lambda)\subseteq\operatorname{\mathrm{loc}}(\cup_{\lambda}\mathbf{p}_{\lambda})\subseteq\operatorname{\mathrm{loc}}(P).$
Thus $\operatorname{\mathrm{loc}}(P)=\operatorname{\mathrm{loc}}(\cup_{\lambda}\mathbf{p}_{\lambda})$. We already saw $\cup_{\lambda}\mathbf{p}_{\lambda}\subseteq\operatorname{thick}(P)$ and so $P$ is proxy-small. ∎
However, closure under cones is rather unusual.
Lemma 6.2.
Proxy-smallness is preserved under taking cones if and only if every object is proxy-small. In particular, this implies every localizing subcategory of $\mathsf{K}$ is generated by objects of $\mathsf{K}^{\mathrm{c}}$.
Proof.
Since the full subcategory of proxy-small objects is closed under suspensions and arbitrary coproducts, it is closed under cones precisely if it is a localizing subcategory. Since the objects of $\mathsf{K}^{\mathrm{c}}$ are proxy-small this occurs if only if every object is proxy-small.
Suppose then that this is the case. If $\mathsf{J}$ is a localizing subcategory then we have
$\mathsf{J}=\operatorname{\mathrm{loc}}(\operatorname{\mathrm{loc}}(X)\mid X\in\mathsf{J})=\operatorname{\mathrm{loc}}(\operatorname{\mathrm{loc}}(X)\cap\mathsf{K}^{\mathrm{c}}\mid X\in\mathsf{J})=\operatorname{\mathrm{loc}}(\mathsf{J}\cap\mathsf{K}^{\mathrm{c}}),$
i.e. $\mathsf{J}$ is generated by the compact objects it contains. ∎
Example 6.3.
Consider $\mathsf{K}=\mathsf{D}_{\{p\}}(\mathbb{Z})=\operatorname{\mathrm{loc}}(\mathbb{F}_{p})\subseteq\mathsf{D}(\mathbb{Z})$. We claim that every object of $\mathsf{K}$ is proxy-small. Indeed, if $X\neq 0$ is an object of $\mathsf{K}$ then there is a triangle
${X}$${X}$${X\otimes\mathbb{F}_{p}}$${\Sigma X}$$\scriptstyle{p}$
where $X\otimes\mathbb{F}_{p}$ is a non-zero sum of suspensions of $\mathbb{F}_{p}$. Thus $\mathbb{F}_{p}\in\operatorname{thick}(X)$ and of course $X\in\operatorname{\mathrm{loc}}(\mathbb{F}_{p})=\mathsf{K}$.
Example 6.4.
The above example generalizes to any non-singular point on a noetherian scheme, i.e. if $(R,\mathfrak{m},k)$ is a regular local ring then every object of $\mathsf{D}_{\{\mathfrak{m}\}}(R)=\operatorname{\mathrm{loc}}(k)$ is proxy-small. One repeats the same argument using that the Koszul complex $K(\mathfrak{m})$ is isomorphic to $k$ and that $X\otimes K(\mathfrak{m})\in\operatorname{thick}(X)$ is both non-zero (if $X\neq 0$) and a sum of suspensions of $k$.
Example 6.5.
For a local noetherian ring $(R,\mathfrak{m},k)$ it need not be the case that every object of $\operatorname{\mathrm{loc}}(k)$ is proxy-small. Indeed, if $R$ is artinian then $\operatorname{\mathrm{loc}}(k)=\mathsf{D}(R)$. If $R$ is not complete intersection then by [16] there is some $X\in\mathsf{D}^{\mathrm{b}}(R)$ which is not proxy-small.
In general, the proxy-small objects are also not closed under taking summands.
Example 6.6.
Consider $\mathsf{K}=\mathsf{D}(\mathbb{Z})$. The object $\mathbb{Q}$ is not proxy-small:
$\operatorname{thick}(\mathbb{Q})=\{\oplus_{i=1}^{m}\Sigma^{a_{i}}\mathbb{Q}^{b_{i}}\mid a_{i}\in\mathbb{Z},b_{i}\in\mathbb{N}\}$
has no non-zero compact object. However $\mathbb{Z}\oplus\mathbb{Q}$ is proxy-small with proxy $\mathbb{Z}$.
Since the proxy-small condition depends only on the thick subcategory generated by an object, it is natural to consider proxy-small thick subcategories, as in Remark 2.1.
Proposition 6.7.
The collection of proxy-small thick subcategories in $\mathsf{K}$, ordered by inclusion, has all least upper bounds.
Proof.
By the same argument as for Lemma 6.1, if $\{P_{\lambda}\mid\lambda\in\Lambda\}$ is a family of proxy-small thick subcategories, then $\operatorname{thick}(P_{\lambda}\mid\lambda\in\Lambda)$ is proxy-small as well. ∎
7. A perspective on complete intersections
We would like to use this new conceptual understanding of proxy-smallness to our benefit. In this section we provide some first applications, present some more examples, and pose some problems.
One setting in which proxy-smallness has been studied extensively is $\mathsf{K}=\mathsf{D}(R)$ for $R$ a commutative noetherian ring. In this case $\mathsf{D}(R)$ satisfies the telescope conjecture by [14] and so Theorem 4.3 applies in full force. This lets us reformulate various results in terms of the coproduct preservation of certain functors.
As an example, let $\phi\colon R\longrightarrow S$ be a map of commutative noetherian rings such that $\phi$ is flat and essentially of finite type. Set
$\mathrm{HH}^{*}(S/R,-)\coloneqq\operatorname{\mathbf{R}Hom}_{S\otimes_{R}S}(S,-)\colon\mathsf{D}(S\otimes_{R}S)\longrightarrow\mathsf{D}(S)\,,$
the Hochschild cohomology functor. Viewing $\mathrm{HH}^{*}(S/R)=\mathrm{HH}^{*}(S/R,S)$ in the enhanced sense, i.e. as a dg algebra or ring spectrum, and combining Theorem 4.3 with the work of Briggs, Iyengar, Letz, and Pollitz [2] we deduce the following corollary to their work.
Corollary 7.1.
With notation as above, $\phi$ is locally complete intersection if and only
$\mathrm{HH}^{*}(S/R,-)\otimes_{\mathrm{HH}^{*}(S/R)}S\colon\mathsf{D}(S\otimes_{R}S)\longrightarrow\mathsf{D}(S)$
preserves coproducts. Moreover, if this is the case then the above gives local cohomology on $\operatorname{Spec}S\times_{\operatorname{Spec}R}\operatorname{Spec}S$ with support on the diagonal $\Delta$ and
$-\otimes_{\mathrm{HH}^{*}(S)}S\colon\mathsf{D}(\mathrm{HH}^{*}(S/R))\longrightarrow\mathsf{D}_{\Delta}(S\otimes_{R}S)$
is a localization. ∎
There are other settings where proxy-smallness of the diagonal appears, although this point of view does not seem to have been made explicit in the literature.
Corollary 7.2.
Let $G$ be a $p$-group and $k$ a field. Then $\mathsf{D}(k(G\times G))$ is a localization of $\mathsf{D}(\mathrm{HH}^{*}(kG))$.
Proof.
Since $k(G\times G)\cong kG^{\mathrm{e}}$ is in $\operatorname{thick}(k)$, it follows that $kG\otimes_{k}k(G\times G)$ is in $\operatorname{thick}(kG)$, which exhibits a non-zero compact object in $\operatorname{thick}(kG)$. Thus $kG$ is proxy-small over $k(G\times G)$. This proves the corollary. ∎
Question 7.3.
Is there a natural description of the kernel of the localization functors $\mathsf{D}(\mathrm{HH}^{*}(S))\longrightarrow\mathsf{D}_{\Delta}(S\otimes_{R}S)$ and $\mathsf{D}(\mathrm{HH}^{*}(kG))\longrightarrow\mathsf{D}(k(G\times G))$?
In fact, it seems in examples that proxy-smallness of the diagonal bimodule and finite generation of Hochschild cohomology occur together. This prompts a definition (which does not seem to have appeared in the literature yet) and some questions.
Definition 7.4.
Let $E$ be a commutative ring spectrum and let $R$ be a commutative $E$-algebra. We say that $R$ is complete intersection (relative to $E$) if $\Delta$ is proxy-small in $\mathsf{D}(R^{\mathrm{op}}\otimes_{E}R)$ where $\Delta$ is $R$ viewed with the obvious bimodule structure (i.e. the bimodule corresponding to the identity functor).
Remark 7.5.
One can modify the above definition by only asking for some relative version of proxy-smallness. For instance, $\mathsf{D}(R^{\mathrm{op}}\otimes_{E}R)$ is monoidal and we could ask for $\otimes$-proxy-smallness of the diagonal. However, since the diagonal is the identity for the monoidal structure this turns out to be trivial.
A good definition should involve some more restrictive closure condition, for instance proxy-smallness relative to the invertible bimodules as in Example 7.9.
Question 7.6.
Greenlees has given several versions of the complete intersection property for ring spectra [9]*Definition 14.3. How does this definition fit into his hierarchy?
Question 7.7.
If the graded-commutative ring $\mathrm{HH}^{*}(R/S)$ is finitely generated is $\Delta$ proxy-small? If $\Delta$ is proxy-small does it imply $\mathrm{HH}^{*}(R/S)$ is finitely generated?
Example 7.8.
It follows from work of Dotsenko, Gélinas and Tamaroff that noncommutative Gorenstein monomial algebras are complete intersection in the sense of Definition 7.4.
Let $Q$ be a finite quiver, $k$ a field, and $I$ an ideal in $kQ$ generated by paths. Assume that the monomial algebra $\Lambda=kQ/I$ is Iwanaga-Gorenstein. According to [4, Theorem 6.3] there is an explicit map $\chi\colon\Sigma^{-p}\Lambda\longrightarrow\Lambda$ in $\mathsf{D}(\Lambda^{\mathrm{op}}\otimes_{k}\Lambda)$ such that, for all finite dimensional $\Lambda$-modules $M$ and $N$, $\operatorname{Ext}^{*}_{\Lambda}(M,N)=\mathrm{HH}^{*}(\Lambda,\operatorname{Hom}_{k}(M,N))$ is finitely generated over the subring $k[\chi]\subseteq\mathrm{HH}^{*}(\Lambda)$; that is, $\Lambda$ satisfies the so-called ${\bf Fg}$ condition. Varying $M$ and $N$ over the simple $\Lambda$-modules we obtain all simple $\Lambda$-bimodules as $\operatorname{Hom}_{k}(M,N)$. Therefore, for each simple bimodule $s$, applying the functor $\operatorname{Ext}_{\Lambda^{\mathrm{op}}\otimes_{k}\Lambda}(-,s)$ to the triangle
[…]
shows that $\operatorname{Ext}_{\Lambda^{\mathrm{op}}\otimes_{k}\Lambda}(\mathrm{cone}(\chi),s)$ is finite dimensional. We can conclude that $\mathrm{cone}(\chi)$ is perfect over $\Lambda^{\mathrm{op}}\otimes_{k}\Lambda$. On the other hand, $\mathrm{cone}(\chi^{n})$ is in $\operatorname{thick}(\mathrm{cone}(\chi))$ for all $n$, and we may recover the diagonal bimodule as the homotopy limit $\Lambda=\operatorname{holim}(\mathrm{cone}(\chi^{n}))$. This shows that $\Lambda\in\operatorname{\mathrm{loc}}(\mathrm{cone}(\chi))$, and so $\Lambda$ is proxy-small in $\mathsf{D}(\Lambda^{\mathrm{op}}\otimes_{k}\Lambda)$ with proxy $\mathrm{cone}(\chi)$.
Likewise, we can deduce that every object $M$ of $\mathsf{D}^{\mathrm{b}}(\Lambda)$ is proxy-small in $\mathsf{D}(\Lambda)$ with proxy $M\otimes_{\Lambda}\mathrm{cone}(\chi)$.
This homological behaviour is similar to that of hypersurface rings in commutative algebra, where $\chi$ is the cohomology operator of Gulliksen. In that case one can always take $p=2$, reflecting the $2$-periodicity of matrix factorizations over hypersurface rings.
The class of Gorenstein monomial algebras includes all gentle algebras by [8].
In other complete-intersection-like examples the diagonal bimodule is not enough, but rather the set of invertible bimodules is proxy-small in the sense of Remark 2.1.
Example 7.9.
Let $q$ be a unit in a field $k$, and consider the quantum complete intersection
$\Lambda=k\langle x,y\rangle/(x^{m},y^{n},xy-qyx).$
For any automorphism $\sigma$ of $\Lambda$ there is an invertible bimodule $\Lambda_{\sigma}$ with underlying $k$-module $\Lambda$ and with left and right action $a\cdot m\cdot c=am\sigma(c)$ (multiplication in $\Lambda$). We will use the two commuting automorphisms $\alpha$ and $\beta$ given by $\alpha(x)=q^{-1}x$, $\alpha(y)=y$, and $\beta(y)=qy$, $\beta(x)=x$ below.
The extension algebra $\operatorname{Ext}^{*}_{\Lambda}(k,k)$ of $\Lambda$ is finite over a subalgebra $k\langle u,v\rangle/(uv-q^{mn}vu)$ [1, 5.3]. Moreover one can compute using [3, 3.1] that, under the map
${\rm HH}^{*}(\Lambda,\Lambda_{\beta^{m}})\longrightarrow{\rm HH}^{*}(\Lambda,k)=\operatorname{Ext}^{*}_{\Lambda}(k,k),$
the class $u$ lifts canonically to a class $\chi_{u}$ in ${\rm HH}^{*}(\Lambda,\Lambda_{\beta^{m}})$. Similarly, $v$ lifts to a class $\chi_{v}$ in ${\rm HH}^{*}(\Lambda,\Lambda_{\alpha^{n}})$. These cohomology operators commute up to a factor of $q^{mn}$, that is, they fit into a commuting square
${\Sigma^{-4}\Lambda}$${\Sigma^{-2}2\Lambda_{\beta^{m}}}$${\Sigma^{-2}2\Lambda_{\alpha^{n}}}$${\Lambda_{\alpha^{n}\beta^{m}}}$$\scriptstyle{\chi_{u}}$$\scriptstyle{\chi_{v}}$$\scriptstyle{\chi_{v}}$$\scriptstyle{q^{mn}\chi_{u}}$
in $\mathsf{D}(\Lambda^{\mathrm{op}}\otimes_{k}\Lambda)$. Finally, we take the colimit of this diagram (that is, take the cone twice) to obtain a quantum Koszul complex $K(\chi_{u},\chi_{v})$ [13]. A similar argument to Example 7.8 shows that $K(\chi_{u},\chi_{v})$ (and all of its twists $K(\chi_{u},\chi_{v})_{\alpha^{i}\beta^{j}}$) is perfect in $\mathsf{D}(\Lambda^{\mathrm{op}}\otimes_{k}\Lambda)$. Just as before, we can also recover $\Lambda$ (and all of its twists $\Lambda_{\alpha^{i}\beta^{j}}$) from the objects $K(\chi_{u},\chi_{v})_{\alpha^{i}\beta^{j}}$ by taking a homotopy limit. Altogether, we have sketched a proof that the set of invertible bimodules $\{\Lambda_{\alpha^{i}\beta^{j}}\mid i,j\in\mathbb{Z}\}$ is proxy-small, with a set $\{K(\chi_{u},\chi_{v})_{\alpha^{i}\beta^{j}}\mid i,j\in\mathbb{Z}\}$ of proxies.
This example is in the spirit of Section 5: in order for the thick closure to be large enough we have closed up with respect to the action of the derived Picard group (or just the subgroup generated by $\Lambda_{\alpha}$ and $\Lambda_{\beta}$, in this case). One can formulate it instead by saying that the diagonal bimodule $\Lambda$ is $\otimes$-proxy-small with respect to the action of the ind-completion of $\operatorname{thick}(\Lambda_{\alpha^{i}\beta^{j}}\mid i,j\in\mathbb{Z})$.
Example 7.10.
A simply connected finite CW-complex $X$ is called rationally elliptic if $\pi_{n}(X)\otimes_{\mathbb{Z}}\mathbb{Q}=0$ for all but finitely many $n$. The results of [10] show that if $X$ is rationally elliptic then the dg algebra $C^{*}(X;\mathbb{Q})$ of rational cochains on $X$ has a proxy-small diagonal, and therefore is complete intersection over $\mathbb{Q}$ in the sense of 7.4.
Rationally elliptic spaces are “spherically complete intersection” in the sense of [10]—they can be constructed, up to rational homotopy equivalence, by starting with a point and using finitely many spherical fibrations. Therefore by [10, Theorem 10.1] they are “endomorphism complete intersection”, that is, there is a sequence of triangles of dg $C^{*}(X\times X;\mathbb{Q})$-modules
$C^{*}(X)=M_{0}\xrightarrow{\ g_{1}\ }\Sigma^{n_{1}}M_{0}\longrightarrow M_{1},\quad\cdots\quad M_{c-1}\xrightarrow{\ g_{c}\ }\Sigma^{n_{c}}M_{c-1}\longrightarrow M_{c}$
such that $M_{c}$ is a perfect dg $C^{*}(X\otimes X;\mathbb{Q})$-module. Each $M_{i}$ has finite dimensional total homology, and by [10, Remark 10.2] each $g_{i}$ has strictly positive cohomological degree. Therefore, as in Example 7.8, we can recover each $M_{i-1}$ from $M_{i}$ by taking a countable homotopy colimit. It follows that $C^{*}(X)$ is proxy-small as a dg module over $C^{*}(X\times X)\simeq C^{*}(X)\otimes C^{*}(X)$ with proxy $M_{c}$, as claimed.
Various other complete-intersection-like conditions on a space are considered in [10]. We also note that the results of [10] are stated much more generally for strongly noetherian rational spaces (we have restricted to finite CW complexes only for the sake of familiarity).
Proxy-smallness does not come cheap
It is natural to ask if there are there are some easy conditions we could put on $P$ and $E=\operatorname{\mathbf{R}Hom}(P,P)$ that would guarantee $P$ is proxy-small. The following example indicates that even imposing some very strong finiteness conditions on $E$ is not sufficient.
Example 7.11.
As we saw in Example 6.6 in $\mathsf{D}(\mathbb{Z})$ the object $\mathbb{Q}$ is not proxy-small, and so having endomorphism ring a field is not enough.
Example 7.12.
Consider $R=k[x,y]/(x,y)^{2}$, for some field $k$, which is an artinian non-Gorenstein ring. We take for $P$ the indecomposable injective module $\operatorname{Hom}_{k}(R,k)$. Then $\operatorname{\mathbf{R}Hom}_{R}(P,P)$ is just $R$. In particular, it is discrete (i.e. no higher homotopical information required) and artinian.
Nevertheless, $P$ is not proxy-small. Indeed, $\operatorname{thick}(P)$ consists of the bounded complexes of injectives with finite dimensional cohomology and so $\operatorname{thick}(P)\cap\mathsf{D}^{\mathrm{perf}}(R)$ consists of complexes with both finite projective and injective dimension. The only such complex can be 00, because $R$ is not Gorenstein; see for example [7, 2.10].
None of this is too surprising. Indeed, proxy-smallness of $P$ depends crucially on how $P$ sits in $\mathsf{K}$ and so it is quite a subtle property. A related question is:
Question 7.13.
What do proxy-small objects form? That is to say, if we fix a compactly generated localizing subcategory $\mathsf{J}$ what is the right structure on the collection of proxy-small objects which generate $\mathsf{J}$?